Fia:

A large fish tank already has 12 gallons of water in it. A hose continues to fill it at a rate of 1/3 of a gallon per minute. Another fish tank already contains 3 gallons of water and is filled at the rate of 1 1/2 times the rate of the first hose. Both hoses are turned on at the same time. Determine the number of minutes until both fish tanks contain the same number of gallons of water.

Ultrilliam:

First things first, let's convert 1/2 and 1/3 to the LCD (Least Common Denominator)

Ultrilliam:

What would the LCD be?

Fia:

The LCD of what...?

Ultrilliam:

1/2 and 1/3

Fia:

6

Ultrilliam:

the LCD is 6, correct. so that means we have 3/6 and 2/6 this will make it easier to calculate how much faster it fills. for every 3/6ths of a gallon, the other tank is filled 1 and 2/6ths times more, what is 1 and 2/6ths times 3/6?

Fia:

2/3

Ultrilliam:

I got this backwards.... 1 and 3/6ths * 2/6ths sorry about that x.x

Fia:

I'm not sure, to be honest ;\

Ultrilliam:

1/2

Ultrilliam:

Which means we have gone full circle... :|... It's been a while since i've done this and I'm very rusty

Ultrilliam:

I don't know where to go from here >_<... lemme get someone who might know

Fia:

Ah, okay. Well, thanks ^^

Hero:

Let x = gallons per minute Let y = number of minutes until both tanks have the same amount of gallons 12 + (1/3)x = y 3 + (1 + 1/2)(1/3)x = y Set y = y, then solve for x

Fia:

^I take it that that's a continuation to what Ultrilliam said?

Hero:

So to sum it up here's a sentence description of what was supposed to be done: Gallons Currently in Tank 1 PLUS Rate of 1st hose TIMES x = y Gallons Currently in Tank 2 PLUS Rate of 2nd hose TIMES rate of 1st hose TIMES x = y

Hero:

Hope that's not too confusing for you.

Fia:

I think I get it...thanks.

Hero:

Great...You're welcome.

Hero:

Let us know what you get for the answer.

Hero:

Actually x = minutes until both tanks have same amount of gallons and y = total gallons of water in both tanks (which should be the same amount).

Hero:

I notice there's still three people currently looking at this question. Has anyone come up with an answer or is everyone still in a state of confusion?

Fia:

Yeah..I'm still confused..

Hero:

I'm sorry, that's my fault. You may have been confused by my initial descriptions for x and y. The truth is, I "rushed" those descriptions because I was pre-occupied with something else I was doing before I saw this question and I was trying to type the solution quickly and get back to what I was doing before. If I went over the steps of how to solve the problem here, do you think that would help you understand better?

Fia:

Yes, please.

Hero:

Okay, here's the full solution for this problem: First, we should let x = minutes until both tanks have same amount of gallons and y = total gallons of water in either tank (which should be the same amount). Next, we should write out the following verbal expression as follows: Gallons Currently in Tank 1 PLUS Rate of 1st hose TIMES x = y Gallons Currently in Tank 2 PLUS Rate of 2nd hose TIMES rate of 1st hose TIMES x = y Afterwards, we input the numbers into the verbal expression: 12 + (1/3)x = y 3 + (1 + 1/2)(1/3)x = y Next set the two expressions equal to each other since y = y 12 + (1/3)x = 3 + (1 + 1/2)(1/3)x Finally, Solve for x: 12 + x/3 = 3 + (1 + 1/2)(x/3) 12 + x/3 = 3 + x/3 + x/6 12 = 3 + x/6 9 = x/6 54 = x So it will take 54 minutes for the tanks to be filled with the same amount of gallons. By the way, if you're curious, 30 gallons is the amount that will be in both tanks after 54 minutes.

Hero:

If anyone still has more questions after this, I can provide a clearer detailed description of the solution steps.